OpenStudy (anonymous):

How do I calculate the cube roots of 8 in the formula a+bj

OpenStudy (anonymous):

8 = 2*2*2

OpenStudy (anonymous):

a + ib ------ 2 + 0i

OpenStudy (anonymous):

use the following formula with n=3 and r=8 use k=0,1,2 \[\Large =r^{\frac{1}{n}}[\cos(\frac{\theta}{n}+\frac{2 \pi k}{n}+isin(\frac{\theta}{n}+\frac{2 \pi k}{n}]\] where \[\theta =0\]

OpenStudy (anonymous):

what ^

OpenStudy (anonymous):

Thanks

hartnn (hartnn):

@bmtana ,did u get it or still have doubts....we can help u better if u post the entire question....

OpenStudy (anonymous):

Here is the question.....get the cube roots of 8 using the formula a+bj and give the main root is exponential form.

OpenStudy (anonymous):

2^3 is the exponential form, and i believe yahoo had it right with his 2 + 0i for the formula a+bi

OpenStudy (anonymous):

Thanks, that helped me a lot.